# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v4

@@ 3. Classification of orbits
Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ with $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
Proof. Assume the contrary. Let $\lambda\neq\lambda^{\prime}$ but $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$. Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$. Then $\gamma$ belongs to the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$. Since both forms lie in $S$, the root $\gamma$ satisfies condition (1). Set $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$. By Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, there exist rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
@@
Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$. Therefore, we may choose our functions $F_{D,\gamma}$ for different $D$ and $\gamma$ in such a way that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$. Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary element of their common orbit, where $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$. Then
@@
[…]
Thus, distinct forms in $S$ have distinct orbits. Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$, then $N.\lambda\neq N.\lambda^{\prime}$.
@@
for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, a straightforward calculation gives
@@
Since each form in $S$ determines its own orbit, the set $S$ accounts for $|S|$ orbits. However, in the work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is computed by group-theoretic methods, and it coincides exactly with $|S|$. Hence, over $\mathbb{F}_{q}$ there are no other orbits. This completes the proof of the main result.
@@
$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ satisfying the following condition: every $\gamma\in\textup{supp}(\lambda)$ fulfills condition (1). We can now state a conjecture.
Conjecture 3.8. For every coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on that orbit.
Matvey A. Surkov: Samara National Research University, Ak. Pavlova 1, Samara, 443011, Russia; Saint Petersburg University, 7/9 Universitetskaya nab., St. Petersburg, 199034, Russia E-mail adress: surkovmatveya@yandex.ru
@@
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3},~\alpha_{4}$
@@
                if(roots.contains(a[i][j])||roots.contains(a[j][i])) { c[i][j]=(m[i][j]-m[j][i])*lambda[a[i][j]+a[j][i]];}
@@
    int x; int[][]a; int[][]m; String[]b; double[][]c; double[]lambda; ArrayList<Integer> roots; ArrayList<Integer> depend;
    boolean q; boolean p; int dep; int last; boolean rank; int n; int count; int rowold; int colold; int rownew; int colnew;

### v5

@@ 3 Classification of orbits
Theorem 3.9.
Let $\lambda,\lambda^{\prime}\in S$ with $\lambda\neq\lambda^{\prime}$. Then $\overline{N}\cdot\lambda\neq\overline{N}\cdot\lambda^{\prime}$.
Proof 3.10.
Assume the contrary. Let $\lambda\neq\lambda^{\prime}$ but $\overline{N}\cdot\lambda=\overline{N}\cdot\lambda^{\prime}$. Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$. Then $\gamma$ belongs to the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$. Since both forms lie in $S$, the root $\gamma$ satisfies condition (1). Set $\mu=\exp(\chi)\cdot\lambda,~\mu^{\prime}=\exp(\chi)\cdot\lambda^{\prime}$. By Theorem 3.2, there exist rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
@@
Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$ and $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ . Therefore, we may choose our functions $F_{D,\gamma}$ for different $D$ and $\gamma$ in such a way that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$. Let $\widetilde{\mu}=\exp(\widetilde{x})\cdot\lambda$ be an arbitrary element of their common orbit, where $\widetilde{x}=\sum_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$. Then
@@
Note that $F_{D,\gamma}(\widetilde{\mu}(e_{\beta}),\beta\succ\gamma)$ does not become infinite because $\widetilde{\mu}(e_{\gamma})$, $\lambda(e_{\gamma})$, and $\lambda^{\prime}(e_{\gamma})$ are well-defined numbers. Hence, $\lambda(e_{\gamma})=\lambda^{\prime}(e_{\gamma})$. This is a contradiction.
Thus, distinct forms in $S$ have distinct orbits. Note that $N\cdot\lambda\subseteq\overline{N}\cdot\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$, then $N\cdot\lambda\neq N\cdot\lambda^{\prime}$.
@@
for the supports 30, 31, 32, see page 2. Thus, a straightforward calculation gives
@@
Since each form in $S$ determines its own orbit, the set $S$ accounts for $|S|$ orbits. However, in the work [3], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is computed by group-theoretic methods, and it coincides exactly with $|S|$. Hence, over $\mathbb{F}_{q}$ there are no other orbits. This completes the proof of the main result.
@@
$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}\cdot\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ satisfying the following condition: every $\gamma\in\operatorname{supp}(\lambda)$ fulfills condition (1). We can now state a conjecture.
Conjecture 3.11.
For every coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on that orbit.
Acknowledgments
This work was supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075–15–2022–287). I thank M. V. Ignatev and A. N. Panov for helpful comments and support during the writing of this paper.
@@ Acknowledgments
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},\alpha_{2},\alpha_{3},\alpha_{4}$
@@
                if(roots.contains(a[i][j])||roots.contains(a[j][i]))
                { c[i][j]=(m[i][j]-m[j][i])*lambda[a[i][j]+a[j][i]];}
@@
    int x; int[][]a; int[][]m; String[]b; double[][]c; double[]lambda;
    ArrayList<Integer> roots; ArrayList<Integer> depend;
    boolean q; boolean p; int dep; int last; boolean rank; int n; int count; int rowold; int colold;
    int rownew; int colnew;